A 4.0 kg block is put on top of a 5.0 kg block. ThFriction of one block on top of another. Full explanation please.
Answer #1

Here there is no relative motion between the blocks till their accelerations are equal.

Draw the free-body diagram of the 5 kg block
mg acts downwards
N acts upwards
Applied force F acts towards right
Frictional force f acts towards left

F - f = Ma
The frictional force is maximum = umg
So
F = Ma + umg = 5a + 0.33*4*9.8 = 5a + 12.93........................(1)

Now draw the free-body diagram for the 4 kg block.
mg acts downwards
N acts upwards
f acts towards right

So f = ma
or umg = ma
or a = ug = 0.33 * 9.8 = 3.234

Substitute in (1) :
F = 5(3.234) + 12.93 = 29.1 N

F= 29.1 N